Use only one variable. Find three consecutive integers such that three times the sum of the first and third integer is nine less than five times the second integer. Thanks
Answer by checkley75(3666)

x+(x+1)+(x+2) are the three numbers.
3[x+(x+2)]=5(x+1)-9
3[2x+2]=5x+5-9
6x+6=5x-4
6x-5x=-4-6
x=-10 answer for the first interger.
-10+1=-9 for the second & -10+2=-8 for the third interger.
proof
3(-10-8)=5*-9-9
3*-18=-45-9
-54=-54